# On the Performance of the Neyman Allocation with Small Pilots

### v3

@@ 4.1 Review on Properties of the FNA
Proposition 4.1 ([24], Theorem 1).
þ Under þLABEL:asm--2nd-moments,asm-1stwave-observation-process,asm-2ndwave-outcomes-device, as $m,n\to\infty$
@@
The above result obtains by applying Theorem 1 in [24, ] to the case with only one stratum. In words, $\widehat{\theta}_{\widetilde{p}}$ and $\widehat{\theta}_{p_{\ast}}$ have the same limiting distribution after suitable centering and scaling. This occurs because in the limit, noise coming from estimating $p_{*}$ with $\widetilde{p}$ is negligible in comparison to the sampling error of the difference-in-means estimator. Researchers employing the large $m$ framework essentially assume that $\widetilde{p}$ is an arbitrarily good estimator for $p_{*}$ so that its sampling error can be ignored. In practice, error in $\widetilde{p}$ can be large, particularly when $m$ is small. Asymptotic approximations that do not take this into account will likely perform poorly in finite sample.
@@
We note two minor differences in the set-up of [24, ]: (1) Their difference-in-means estimator uses both main wave and pilot observations. The main wave allocation therefore needs to be adjusted since half of the pilot is treated. (2) They assume that $m$ is of the same order as $n$. These differences do not affect the result.
@@
To better understand the behavior of the FNA under small pilots, we follow [18, ] in studying its properties in an asymptotic framework that takes $m$ to be fixed, even as $n\to\infty$.
A growing literature in econometrics uses fixed-sample asymptotics to study settings in which the “effective sample size” is small. For example, when data exhibit cluster-dependence, settings with few clusters pose unique challenges for estimation and inference. To tailor their analyses to these problems, papers such as [27, ], [13, ] and [14, ] employ asymptotic frameworks in which the number of clusters is fixed in the limit. Similarly, to model difference-in-differences studies involving few treated units, [17, ] keep the number of treated units fixed even as the number of untreated units tend to infinity. In the same vein, inference and specification testing for regression discontinuity designs typically involves few observations around the discontinuity. To capture this, [12, ] analyze a permutation test for continuity in the distribution of baseline covariates under an asymptotic regime with a fixed number of observations on either side of the discontinuity.
The small pilot approach is similar in spirit to these papers. In keeping $m$ fixed, $\widetilde{p}$ is a noisy estimate of $p_{*}$ even in the limit. Preserving this important feature of the statistical problem makes our chosen framework more appropriate for analyzing experiments with small pilots. In this setting, $\widehat{\theta}_{\widetilde{p}}$ converges in distribution to a mixture of Gaussians instead of $\mathcal{N}(0,\Sigma_{*})$. Specifically, the form of the limiting mixture distribution depends on the distribution of $\widetilde{p}$. This is the content of the following result due to [18, ] (also see the discussion following equation (3.5) of [19, ]):
Proposition 4.2 ([18], Theorem 3.17).
þ Under þLABEL:asm--2nd-moments,asm-1stwave-observation-process,asm-2ndwave-outcomes-device, if $m$ remains fixed as $n\to\infty$,
@@
The above proposition can be obtained from Theorem 3.17 of [18, ] by specializing it to the case with unstratified sampling and assignment ($q=1,\psi=1$). Conditional normality then implies a mixed normal limiting distribution by a dominated convergence argument. Also see related intuition below.
@@
Earlier versions of this paper present the same mixed normality result but derived under simple random assignment, which is not directly implied by [18, ].
@@
þ The weak limit $\mathcal{L}_{m}$ in þ4.2 has mean zero. To see this, note that conditional on a value of the assignment probability $p$, i.e. conditional on the event $\widetilde{p}=p$, $\mathcal{L}_{m}$ has a ${N}(0,s(p))$ distribution (where $s(p)$ is as defined in þ4.2). The conclusion then holds by the Law of Iterated Expectations.
@@
$\left\{\sqrt{n}\left(\widehat{\theta}_{p}-\theta\right):p\in(0,1)\right\}\leavevmode\nobreak\ ,$
@@ 4.2 When is the FNA preferred?
Proposition 4.2 shows that when pilots are small, $\hat{\theta}_{\tilde{p}}$ has asymptotic distribution that is different from $\hat{\theta}_{p_{\ast}}$. It stands to reason that the FNA may not be optimal in this regime. In this sub-section we compare the asymptotic variances of $\widehat{\theta}_{\widetilde{p}}$, $\widehat{\theta}_{p_{\ast}}$ as well as $\widehat{\theta}_{p}$ with $p=\frac{1}{2}$. We first show that $\widehat{\theta}_{\widetilde{p}}$ has larger asymptotic variance in the fixed-$m$ regime than in the large-$m$ regime. We next show that under reasonable ranges of parameter values, the asymptotic variance of $\widehat{\theta}_{\widetilde{p}}$ can exceed that of $\widehat{\theta}_{p}$ with $p=\frac{1}{2}$.
@@
þ By þ4.2 and þ4.4, $\widehat{\theta}_{\widetilde{p}}$ has a limit distribution which has mean zero. Given that $\widehat{\theta}_{p_{\ast}}$ and $\widehat{\theta}_{p}$ also have mean zero normal limit distributions, comparing asymptotic mean squared errors is equivalent to comparing the variances of the limit distributions.
@@
Under þLABEL:asm--2nd-moments,asm-1stwave-observation-process,asm-2ndwave-outcomes-device, suppose $m$ remains fixed as $n\to\infty$. $\mathcal{L}_{m}$ has variance:
$\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]>\Sigma_{*}\leavevmode\nobreak\ .$
@@
$Z_{m}=\frac{\widetilde{\sigma}_{m}(1)}{\sigma(1)}\bigg{/}\frac{\widetilde{\sigma}_{m}(0)}{\sigma(0)}\leavevmode\nobreak\ .$
@@
$B_{m}=\frac{1}{2}\,\mathbb{E}\left[\frac{1}{Z_{m}}+Z_{m}\right]\leavevmode\nobreak\ .$
@@
Under þLABEL:asm--2nd-moments,asm-1stwave-observation-process,asm-2ndwave-outcomes-device, suppose $m$ remains fixed as $n\to\infty$. Then
$C_{m}=\left[\frac{1}{c_{m}}\,,\,c_{m}\right]\leavevmode\nobreak\ .$ (4.1)
@@
[…]
@@
The exact properties of $C_{m}$ depends on the underlying distributions of potential outcomes. To understand its behavior in a more general setting, we study its first-order approximation under a sub-Gaussian assumption. Recall that $F$ is the distribution of $(Y(0),Y(1))$ in Assumption 2.1.
@@
$\mathbb{P}\left(|X|\geq t\right)\leq 2\exp\left(-t^{2}/K^{2}\right)\quad\mbox{for all }t\geq 0$
@@
Under þLABEL:asm--2nd-moments,asm-1stwave-observation-process,asm-2ndwave-outcomes-device, suppose $n\to\infty$. Suppose additionally that $F$ is sub-Gaussian. Then,
@@
$V=\frac{1}{4}\left(\frac{\mathbb{E}\left[\left(Y_{i}(1)-\mu(1)\right)^{4}\right]}{\sigma^{4}(1)}+\frac{\mathbb{E}\left[\left(Y_{i}(0)-\mu(0)\right)^{4}\right]}{\sigma^{4}(0)}-2\right)\leavevmode\nobreak\ .$
@@
Furthermore, $|C_{m}|$ is shrinking to 00 at the rate $1/\sqrt{m}$. Letting $m\to\infty$, $\widehat{\theta}_{\widetilde{p}}$ has weakly lower asymptotic variance across all parameter values. Hence, we recover the classic result concerning the optimality of the Neyman Allocation. When $m$ is small, however, $C_{m}$ can be wide, as Proposition 4.4 suggests. As we will argue in Section 5, many empirical applications have $\sigma(1)/\sigma(0)$ close to $1$, so that for small $m$, they fall within the range in which balanced randomization is preferred.
@@ 4.3 Efficiency Loss from the FNA
$\mathcal{L}^{d}(F)=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]-2\left(\sigma^{2}(1)+\sigma^{2}(0)\right)\leavevmode\nobreak\ ,$
$\quad\mathcal{L}^{r}(F)=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]\bigg{/}2\left(\sigma^{2}(1)+\sigma^{2}(0)\right)\leavevmode\nobreak\ .$
@@
Under þLABEL:asm--2nd-moments,asm-1stwave-observation-process,asm-2ndwave-outcomes-device, suppose $m$ remains fixed as $n\to\infty$. Then
$\mathcal{L}^{d}(F)$ $=2\left(B_{m}-1\right)\sigma(1)\sigma(0)-\left(\sigma(1)-\sigma(0)\right)^{2}\leavevmode\nobreak\ ,$
$\mathcal{L}^{r}(F)$ $=\frac{1}{2}+B_{m}\cdot\frac{\sigma(1)\sigma(0)}{\sigma^{2}(1)+\sigma^{2}(0)}\leavevmode\nobreak\ .$
@@
$\mathcal{L}^{d}(F)$ $=\frac{V}{m}\cdot\sigma(1)\sigma(0)-\left(\sigma(1)-\sigma(0)\right)^{2}+\delta^{d}_{m}\leavevmode\nobreak\ ,$
$\mathcal{L}^{r}(F)$ $=\frac{1}{2}+\left(1+\frac{V}{2m}\right)\cdot\frac{\sigma(1)\sigma(0)}{\sigma^{2}(1)+\sigma^{2}(0)}+\delta^{a}_{m}\leavevmode\nobreak\ ,$
where $\delta^{d}_{m},\delta^{a}_{m}=o(1/m)$.
The above proposition shows us how the efficiency loss from using the FNA depends on $B_{m}$ as well as the amount of heteroskedasticity in $F$. To better understand the trade-off from the two terms, suppose we fix $B_{m}$ and reparametrize $\sigma(0)=1/\sqrt{1+\rho^{2}}$, $\sigma(1)=\rho/\sqrt{1+\rho^{2}}$. Now let us consider the effect of changing $\rho$. This is the thought experiment in which we change the relative scaling of $\varepsilon(1)=Y(1)-\mu(1)$ and $\varepsilon(0)=Y(0)-\mu(0)$ while keeping their functional forms fixed. Additionally, this entails keeping $\text{AVar}(\hat{\theta}_{\frac{1}{2}})=\text{Var}(Y(1))+\text{Var}(Y(0))$ fixed at $2$. This scaling is useful for discussing the properties of $\mathcal{L}^{d}$ since it scales linearly with $\sigma(0)$ and $\sigma(1)$. It is irrelevant to $\mathcal{L}^{r}$ which is invariant to the scale of $\sigma(0)$ and $\sigma(1)$. With this parametrization, we have that:
$\frac{\partial\mathcal{L}^{d}}{\partial\rho}=\frac{2B_{m}(1-\rho^{2})}{(1+\rho^{2})^{2}}\quad\mbox{ and }\quad\frac{\partial\mathcal{L}^{a}}{\partial\rho}=\frac{B_{m}\left(1-\rho^{2}\right)}{(1+\rho^{2})^{2}}\leavevmode\nobreak\ .$ (4.2)
The above expressions are intuitive. When $\rho$ exceeds $1$, increasing $\rho$ amounds to increasing heteroskedasticity. FNA is more preferable to balanced randomization and the ratio loss decreases.
Furthermore, it straightforward to see that both $\mathcal{L}^{d}$ and $\mathcal{L}^{r}$ are maximized when $\rho=1$, with the respective values of $B_{m}$ and $\frac{1}{2}(1+B_{m})$. On the other hand, $\mathcal{L}^{d}$ and $\mathcal{L}^{r}$ have minimum values $-1$ and $\frac{1}{2}$ respectively, which are obtained when either $\rho=0$ or $\rho\to\infty$. In summary, for a given functional form of $\varepsilon(0)$ and $\varepsilon(1)$, the loss from using FNA depends on $B_{m}$ but the maximum possible gain does not. This asymmetry foreshadows the following result.
Since $F$ is unknown, we might try to make headway by comparing the supremum and infimum of $\mathcal{L}^{d}(F)$ and $\mathcal{L}^{r}(F)$ (i.e. maximum losses and gains) over appropriate classes of DGPs. To that end, consider the following.
@@
Under þLABEL:asm--2nd-moments,asm-1stwave-observation-process,asm-2ndwave-outcomes-device, suppose $m$ remains fixed as $n\to\infty$. Then,
$\sup_{F\in\mathcal{F}(K)}\mathcal{L}^{d}(F)=\infty\quad$ $,\quad\inf_{F\in\mathcal{F}(K)}\mathcal{L}^{d}(F)=-K$
$\sup_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\infty\quad$ $,\quad\inf_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\frac{1}{2}$
In words, the worst case loss from using the FNA – whether measured as a difference or as a ratio relative to the variance of the balanced allocation – is unbounded. On the other hand, the gains are bounded. Most strikingly, the fact that $\inf_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\frac{1}{2}$ suggests that the best a researcher can do is to obtain main wave confidence intervals for $\hat{\theta}$ that are $1/\sqrt{2}\approx 70\%$ the length of that from balanced randomization. This comes at the cost of potentially doing much worse as a result of noise from the pilot estimation.
We can also interpret the result in terms of regret. Suppose a researcher is constrained to either using the FNA or balanced randomization. If their utility over the second stage estimator is given by $U(\hat{\theta})=-\left({\hat{\theta}-\theta}\right)^{2}$, then $\mathcal{L}^{d}(F)$ is the asymptotic regret of choosing the FNA, while $-\mathcal{L}^{d}(F)$ is the asymptotic regret of choosing balanced randomization. A similar argument can be made for $\mathcal{L}^{r}(F)$ with $U(\hat{\theta})=-\log\left({\hat{\theta}-\theta}\right)$. Proposition 4.7 then says that the regret of using the FNA over balanced randomization is unbounded, while the regret of using balanced randomization over the FNA is not. Balanced randomization is therefore preferred in terms of maximum regret. This is admittedly a very conservative way of choosing an estimation procedure. However, it is coherent in the following sense. If we argue for the use of that FNA on the grounds that there may be a lot of heteroskedasticity in the population, then we are making an argument about the (pointwise) regret of the methods. However, if researchers are being agnostic about $\sigma(1)/\sigma(0)$, they should also be agnostic about $F$. If that is the case, then we should also consider regret over all plausible DGPs.
@@
Proposition 4.7 continues to hold if we restrict $F$ to that $Y(0)$ and $Y(1)$ have uniformly bounded $a^{\text{th}}$ moments for any $a$.
@@
It is also known in the large pilot case that $\inf_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\frac{1}{2}$. See for example [32, ] and [36, ].

### v4

@@ 4.1 Review on Properties of the FNA
Proposition 4.1 (Hahn et al. (2011), Theorem 1).
Under Assumptions 2.1, 2.2 and 2.3, as $m,n\to\infty$
@@
The above result obtains by applying Theorem 1 in Hahn et al. (2011) to the case with only one stratum. In words, $\widehat{\theta}_{\widetilde{p}}$ and $\widehat{\theta}_{p_{\ast}}$ have the same limiting distribution after suitable centering and scaling. This occurs because in the limit, noise coming from estimating $p_{*}$ with $\widetilde{p}$ is negligible in comparison to the sampling error of the difference-in-means estimator. Researchers employing the large $m$ framework essentially assume that $\widetilde{p}$ is an arbitrarily good estimator for $p_{*}$ so that its sampling error can be ignored. In practice, error in $\widetilde{p}$ can be large, particularly when $m$ is small. Asymptotic approximations that do not take this into account will likely perform poorly in finite sample.
@@
We note two minor differences in the set-up of Hahn et al. (2011): (1) Their difference-in-means estimator uses both main wave and pilot observations. The main wave allocation therefore needs to be adjusted since half of the pilot is treated. (2) They assume that $m$ is of the same order as $n$. These differences do not affect the result.
@@
To better understand the behavior of the FNA under small pilots, we follow Cytrynbaum (2021) in studying its properties in an asymptotic framework that takes $m$ to be fixed, even as $n\to\infty$.
A growing literature in econometrics uses fixed-sample asymptotics to study settings in which the “effective sample size" is small. For example, when data exhibit cluster-dependence, settings with few clusters pose unique challenges for estimation and inference. To tailor their analyses to these problems, papers such as Ibragimov and Müller (2010), Canay et al. (2017) and Canay et al. (2021) employ asymptotic frameworks in which the number of clusters is fixed in the limit. Similarly, to model difference-in-differences studies involving few treated units, Conley and Taber (2011) keep the number of treated units fixed even as the number of untreated units tend to infinity. In the same vein, inference and specification testing for regression discontinuity designs typically involves few observations around the discontinuity. To capture this, Canay and Kamat (2017) analyze a permutation test for continuity in the distribution of baseline covariates under an asymptotic regime with a fixed number of observations on either side of the discontinuity.
The small pilot approach is similar in spirit to these papers. In keeping $m$ fixed, $\widetilde{p}$ is a noisy estimate of $p_{*}$ even in the limit. Preserving this important feature of the statistical problem makes our chosen framework more appropriate for analyzing experiments with small pilots. In this setting, $\widehat{\theta}_{\widetilde{p}}$ converges in distribution to a mixture of Gaussians instead of $\mathcal{N}(0,\Sigma_{*})$. Specifically, the form of the limiting mixture distribution depends on the distribution of $\widetilde{p}$. This is the content of the following result due to Cytrynbaum (2021) (also see the discussion following equation (3.5) of Cytrynbaum (2023)):
Proposition 4.2 (Cytrynbaum (2021), Theorem 3.17).
Under Assumptions 2.1, 2.2 and 2.3, if $m$ remains fixed as $n\to\infty$,
@@
The above proposition can be obtained from Theorem 3.17 of Cytrynbaum (2021) by specializing it to the case with unstratified sampling and assignment ($q=1,\psi=1$). Conditional normality then implies a mixed normal limiting distribution by a dominated convergence argument. Also see the related intuition below.
@@
Earlier versions of this paper present the same mixed normality result but derived under simple random assignment, which is not directly implied by Cytrynbaum (2021).
@@
The weak limit $\mathcal{L}_{m}$ in Proposition 4.2 has mean zero. To see this, note that conditional on a value of the assignment probability $p$, i.e. conditional on the event $\widetilde{p}=p$, $\mathcal{L}_{m}$ has a ${N}(0,s(p))$ distribution (where $s(p)$ is as defined in Proposition 4.2). The conclusion then holds by the Law of Iterated Expectations.
@@
$\left\{\sqrt{n}\left(\widehat{\theta}_{p}-\theta\right):p\in(0,1)\right\}~{},$
@@ 4.2 When is the FNA preferred?
Proposition 4.2 shows that when pilots are small, $\hat{\theta}_{\tilde{p}}$ has asymptotic distribution that is different from $\hat{\theta}_{p_{\ast}}$. It stands to reason that the FNA may not be optimal in this regime. In this sub-section, we compare the asymptotic variances of $\widehat{\theta}_{\widetilde{p}}$, $\widehat{\theta}_{p_{\ast}}$ as well as $\widehat{\theta}_{p}$ with $p=\frac{1}{2}$. We first show that $\widehat{\theta}_{\widetilde{p}}$ has larger asymptotic variance in the fixed-$m$ regime than in the large-$m$ regime. We next show that under reasonable ranges of parameter values, the asymptotic variance of $\widehat{\theta}_{\widetilde{p}}$ can exceed that of $\widehat{\theta}_{p}$ with $p=\frac{1}{2}$.
@@
It can be shown that $\widehat{\theta}_{\widetilde{p}}$, $\widehat{\theta}_{p_{\ast}}$ and $\widehat{\theta}_{p}$ are all unbiased estimators of $\theta$ in finite samples. Therefore, comparing asymptotic mean squared errors is equivalent to comparing the variances of the limit distributions.
@@
Under Assumptions 2.1, 2.2 and 2.3, suppose $m$ remains fixed as $n\to\infty$. $\mathcal{L}_{m}$ has variance:
$\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]>\Sigma_{*}~{}.$
@@
$Z_{m}=\frac{\widetilde{\sigma}_{m}(1)}{\sigma(1)}\bigg{/}\frac{\widetilde{\sigma}_{m}(0)}{\sigma(0)}~{}.$
@@
$B_{m}=\frac{1}{2}\,\mathbb{E}\left[\frac{1}{Z_{m}}+Z_{m}\right]~{}.$
@@
Under Assumptions 2.1, 2.2 and 2.3, suppose $m$ remains fixed as $n\to\infty$. Then
$C_{m}=\left[\frac{1}{c_{m}}\,,\,c_{m}\right]~{}.$
@@
$c_{m}=B_{m}+\sqrt{B_{m}^{2}-1}>1\quad\text{and}\quad\left\lvert C_{m}\right\rvert=2\sqrt{B_{m}^{2}-1}\;>\;0~{}.$
@@
The exact properties of $C_{m}$ depend on the underlying distributions of potential outcomes. To understand its behavior in a more general setting, we study its first-order approximation under a sub-Gaussian assumption. Recall that $F$ is the distribution of $(Y(0),Y(1))$ in Assumption 2.1.
@@
$\mathbb{P}\left(|X|\geq t\right)\leq 2\exp\left(-t^{2}/K^{2}\right)\quad\mbox{for all }t\geq 0~{}.$
@@
Under Assumptions 2.1, 2.2 and 2.3, suppose $n\to\infty$. Suppose additionally that $F$ is sub-Gaussian. Then,
@@
$V=\frac{1}{4}\left(\frac{\mathbb{E}\left[\left(Y_{i}(1)-\mu(1)\right)^{4}\right]}{\sigma^{4}(1)}+\frac{\mathbb{E}\left[\left(Y_{i}(0)-\mu(0)\right)^{4}\right]}{\sigma^{4}(0)}-2\right)~{}.$
@@
Furthermore, $|C_{m}|$ is shrinking to 00 at the rate $1/\sqrt{m}$. Letting $m\to\infty$, $\widehat{\theta}_{\widetilde{p}}$ has weakly lower asymptotic variance across all parameter values. Hence, we recover the classic result concerning the optimality of the Neyman Allocation. When $m$ is small, however, $C_{m}$ can be wide, as Proposition 4.4 suggests. As we will argue in Section 5, many empirical applications have $\sigma(1)/\sigma(0)$ close to $1$, so that for small $m$, these ratios fall within the range in which balanced randomization is preferred.
@@ 4.3 Efficiency Loss from the FNA
$\mathcal{L}^{d}(F)=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]-2\left(\sigma^{2}(1)+\sigma^{2}(0)\right)~{},$
$\quad\mathcal{L}^{r}(F)=\mathbb{E}\left[\frac{\sigma^{2}(1)}{\widetilde{p}}+\frac{\sigma^{2}(0)}{1-\widetilde{p}}\right]\bigg{/}2\left(\sigma^{2}(1)+\sigma^{2}(0)\right)~{}.$
@@
Under Assumptions 2.1, 2.2 and 2.3, suppose $m$ remains fixed as $n\to\infty$. Then
$\mathcal{L}^{d}(F)$ $=2\left(B_{m}-1\right)\sigma(1)\sigma(0)-\left(\sigma(1)-\sigma(0)\right)^{2}~{},$
$\mathcal{L}^{r}(F)$ $=\frac{1}{2}+B_{m}\cdot\frac{\sigma(1)\sigma(0)}{\sigma^{2}(1)+\sigma^{2}(0)}~{}.$
@@
$\mathcal{L}^{d}(F)$ $=\frac{V}{m}\cdot\sigma(1)\sigma(0)-\left(\sigma(1)-\sigma(0)\right)^{2}+\delta^{d}_{m}~{},$
$\mathcal{L}^{r}(F)$ $=\frac{1}{2}+\left(1+\frac{V}{2m}\right)\cdot\frac{\sigma(1)\sigma(0)}{\sigma^{2}(1)+\sigma^{2}(0)}+\delta^{r}_{m}~{},$
where $\delta^{d}_{m},\delta^{r}_{m}=o(1/m)$.
The above proposition shows us how the efficiency loss from using the FNA depends on $B_{m}$ as well as the amount of heteroskedasticity in $F$. To better understand the trade-off from the two terms, suppose we fix $B_{m}$ and reparametrize $\sigma(0)=1/\sqrt{1+\rho^{2}}$, $\sigma(1)=\rho/\sqrt{1+\rho^{2}}$. Now let us consider the effect of changing $\rho$. This is the thought experiment in which we change the relative scaling of $\varepsilon(1)=Y(1)-\mu(1)$ and $\varepsilon(0)=Y(0)-\mu(0)$ while keeping their functional forms fixed. Additionally, this entails keeping $\text{AVar}(\hat{\theta}_{\frac{1}{2}})=\text{Var}(Y(1))+\text{Var}(Y(0))$ fixed at $2$. This scaling is useful for discussing the properties of $\mathcal{L}^{d}$ since it scales linearly with $\sigma(0)$ and $\sigma(1)$. $\mathcal{L}^{r}$ is invariant to the scale of $\sigma(0)$ and $\sigma(1)$. With this parametrization, we have that:
$\frac{\partial\mathcal{L}^{d}}{\partial\rho}=\frac{2B_{m}(1-\rho^{2})}{(1+\rho^{2})^{2}}\quad\mbox{ and }\quad\frac{\partial\mathcal{L}^{r}}{\partial\rho}=\frac{B_{m}\left(1-\rho^{2}\right)}{(1+\rho^{2})^{2}}~{}.$
The above expressions are intuitive. When $\rho$ exceeds $1$, increasing $\rho$ amounts to increasing heteroskedasticity. FNA is more preferable to balanced randomization and the ratio loss decreases.
Furthermore, it straightforward to see that both $\mathcal{L}^{d}$ and $\mathcal{L}^{r}$ are maximized when $\rho=1$, with the respective values of $B_{m}$ and $\frac{1}{2}(1+B_{m})$. On the other hand, $\mathcal{L}^{d}$ and $\mathcal{L}^{r}$ have minimum values $-1$ and $\frac{1}{2}$ respectively, which are obtained when either $\rho=0$ or $\rho\to\infty$. In summary, for a given functional form of $\varepsilon(0)$ and $\varepsilon(1)$, the loss from using FNA depends on $B_{m}$ but the maximum possible gain does not. This asymmetry foreshadows the next result.
Since $F$ is unknown, we proceed by comparing the supremum and infimum of $\mathcal{L}^{d}(F)$ and $\mathcal{L}^{r}(F)$ (i.e. maximum losses and gains) over appropriate classes of DGPs. To that end, consider the following.
@@
Under Assumptions 2.1, 2.2 and 2.3, suppose $m$ remains fixed as $n\to\infty$. Then,
$\sup_{F\in\mathcal{F}(K)}\mathcal{L}^{d}(F)=\infty\quad$ $,\quad\inf_{F\in\mathcal{F}(K)}\mathcal{L}^{d}(F)=-K~{},$
$\sup_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\infty\quad$ $,\quad\inf_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\frac{1}{2}~{}.$
In words, the worst case loss from using the FNA – whether measured as a difference or as a ratio relative to the variance of the balanced allocation – is unbounded. On the other hand, the gains are bounded. Most strikingly, the fact that $\inf_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\frac{1}{2}$ suggests that the best a researcher can do with the FNA is to obtain main wave confidence intervals for $\hat{\theta}$ that are $1/\sqrt{2}\approx 70\%$ the length of that from balanced randomization. This comes at the cost of potentially doing much worse as a result of noise from the pilot estimation.
We can also interpret the result in terms of regret. Suppose a researcher is constrained to either using the FNA or the balanced allocation. If their utility over the second stage estimator is given by $U(\hat{\theta})=-\left({\hat{\theta}-\theta}\right)^{2}$, then $\mathcal{L}^{d}(F)$ is the asymptotic regret of choosing the FNA, while $-\mathcal{L}^{d}(F)$ is the asymptotic regret of choosing the balanced allocation. A similar argument can be made for $\mathcal{L}^{r}(F)$ with $U(\hat{\theta})=-\log\left({\hat{\theta}-\theta}\right)$. Proposition 4.7 then says that the regret of using the FNA over the balanced allocation is unbounded, while the regret of using the balanced allocation over the FNA is not. The balanced allocation is therefore preferred in terms of maximum regret. This is admittedly a very conservative way of choosing an estimation procedure. However, it is coherent in the following sense. Researchers may prefer the FNA if they want to guard against large amounts of heteroskedasticity in the population. However, if researchers are being agnostic about $\sigma(1)/\sigma(0)$, they should also be agnostic about $F$, in which case the balanced allocation may ultimately be preferable.
@@
Proposition 4.7 continues to hold if we further restrict $F$ so that $Y(0)$ and $Y(1)$ have uniformly bounded $a^{\text{th}}$ moments for any $a$.
@@
It is also known in the large pilot case that $\inf_{F\in\mathcal{F}(\infty)}\mathcal{L}^{r}(F)=\frac{1}{2}$. See for example Obradovic (2023) and Zhao (2023).